isten:

mrs.yap owns 45 bonds with a oar value of 1,000 each and pays 8 1/2% interest.what annual income does mrs.yap get from these bonds?

jhonyy9:

so like a first step 8 1/2% = (8*2+1)/2 = 17/2 = 8,5% how you think 45 bonds with value of 1000 each will be equal by what sum in total ? after get this sum so just you need calcule the 8,5% of this sum and in this way will get the right answer sure easy @supie your opinion pls about this ? ty